A metallic rod of length L and cross section area A is made of material of Young`s modulus Y. If the rod is elongated by an amount y then work done is proportional to ???

Grade:11

2 Answers

Ishita
7 years ago
Y=F/A       [here F= force applied]
    y/L
 
F= YAy
      L
 
dW=F dy
∫dW= 0y F dy
W= YA (y)2
       2L
W= YAy  X  y
       L          2
 
W= ½ (F) (y)
W is proportional to y
Ishita
7 years ago
 
Y=F/A       [here F= force applied]
    y/L
 
F= YAy
      L
 
dW=F dy
∫dW= 0y F dy
W= YA (y)2
       2L
W= YAy  X  y
       L          2
 
W= ½ (F) (y)
W is proportional to y